From GlueXWiki
Jump to: navigation, search

September 22, 2011 FDC meeting


  1. Production Construction Tracking (Dave)
    • Status
    • First package tests (Lubomir)
  2. Engineering update (Bill)
  3. Electronics update (Chris)
  4. Chamber testing at EEL126 (Beni)
  5. Other: latest pictures of the package


Participants: Eugene, Bill, Dave, Chris, Simon, Beni, Vladimir, Mark, and Lubomir. Vladimir is the new graduate student from MePhi Moscow.


- Dave: started checking the continuity before taking the frame out of the stringing table: on the wire frame for cell#5 (that was just finished) one broken wire was found and was fixed today. Epoxy will cure overnight and tomorrow we will put on the stringing table the wire frame for cell#3 to fix one wire there too. Wire frame for cell#4 is ready, just needs final cleaning. On the last cathode (for the first package) the rigid-flexes were glued to the cathodes. Next, they will be tested with a generator and then the daughter cards glued.

- We discussed the reason for these broken wires, so far we had always one broken wire per plane. On the last one found on the frame for cell#5, we looked with a microscope. The wire was broken at the end of the epoxy right where it touches the soldering pad. The wire tip was bent at ~90 deg at this place somehow. It was not clear if the tip was inside or outside of the epoxy. The wire was checked by doing tension measurements on it. The only operation that happened later was cutting the wires and then doing the continuity measurements. One explanation could be a cold solder and then the wire brakes when twisting it. Dave suggested also that if the wire is above the PCB it may not be fully covered with epoxy and is easily breakable a that point.

- End of last week we put three cells in the package after replacing two wires in the second cell and changing all the O-rings. The first three cells were tested channel by channel by Vladimir. All were OK except no signal on one wire channel in the new third cell, most likely broken wire. The third cell was conditioned with HV relatively fast, but then suddenly developed very high current; we left it over the night at 1kV and then in the morning it was OK again.

- We found again gas leakage with the three cells in about the same area as with the two cells. The amount of leakage was about the same (bubbling above 80-100 ccpm) as with the two cells, indicating that the problem is in the second cell. Bill suspects it is the seam on the wire frame between the PCBs that may have been somehow damaged during the installation. To fix the broken wire and look at the leakage place we decided to unstack the package down to the second wire frame. After finding the broken wire on the frame that is on the stringing table we decided to continue with this tomorrow.


- Bill: the new fixture and hubs with adjustable height should be ready by the end of the month. Probably we will use them for the second package. The first fixture also can be refurbished to be used with the new hubs. Two more gusset rings ready tomorrow. Parts for the HV brackets and cooling system are also out for production. After the meeting: the package support brackets will be at JLab today.


- Anatoly tested several fresh PCBs: no broken resistors, but few resistors down to 50%. So, most likely the resistors were broken by us while making the frames, but the resistors with deviations from the nominal value came from the manufacturer. Fernando was surprised (before the meeting) and suggested replacing those that are more than 10% off, although these are current limiting resistors, i.e. not so critical. No explanation how this may have happened during the population of the boards, since these resistors are at different places on the different boards.

Chamber testing at 126

- Beni raised several issues about the testing set-up. Designed with Mark a frame for the cable trays. Not clear how to organize the low voltage supply, should discuss it with Fernando. Also there's no final design yet of the cooling system; we discussed also what liquid to be used. fADC125: Gerard thinks the last FPGA (that gets the data from all the channels) doesn't have enough power for the needed functions and has to be replaced with a better one (according to Cody). The modifications will delay the production. The hope is that there will be pre-production of 20 units and we can use a good fraction of them for the FDC tests.

- Lubomir: we moved the second cell from 126 to Blue Crab and opened it. The wire plane had ~2 MOhm between field and sense wires on HV sector #4 and that was confirmed after opening. A spot was found close to a field and sense wire at the signal side in the same sector. Also some corrosion is visible on the soldering pads close to the spot, indicating it could be reactive. Trying to clean that spot reflects directly the resistance. Not clear what the material is: semi-transparent like epoxy or Humi-seal?, can be removed by a wooden stick. Not clear at all how did it get there inside the gas volume.


New pictures of the three cell package are linked above.